i want to know a quadratic equation whose roots are real and non congugate


 

Grade:12

5 Answers

sneh gup
10 years ago

x^2-5X+6=0

Vikas TU
10 years ago

x^2 - 3x + 2 = 0

it has real and distinct roots
thus no imaginary and no conjugates! 

Akash Kumar Dutta
10 years ago

infinitely many...
eg. x^2-x+6=0...roots are real and non conjugate!
i think you wanted to know something else....cause it doesnt seems to be any doubt!

Rakshith Gore
10 years ago

xsquare-5x+6

Abhishekh kumar sharma
10 years ago

infinitely many except for equations whose d<0